40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon May 4 10:38:52 2009, steve, Update, MOPA, laser power is dropped dtecup.jpg
    Reply  Mon May 4 16:49:56 2009, Alberto, Update, MOPA, laser power is dropped 
    Reply  Tue May 5 10:42:18 2009, steve, Update, MOPA, laser power is back uncloged.jpg
Message ID: 1543     Entry time: Mon May 4 16:49:56 2009     In reply to: 1542
Author: Alberto 
Type: Update 
Category: MOPA 
Subject: laser power is dropped 

Quote:

As PSL-126MOPA_DTEC went up, the power out put went down yesterday

Alberto, Jenne, Rob, Steve,
 
later on in the afternoon, we realized that the power from the MOPA was not recovering and we decided to hack the chiller's pipe that cools the box.
 
Without unlocking the safety nut on the water valve inside the box, Jenne performed some Voodoo and twisted a bit the screw that opens it with a screw driver. All the sudden some devilish bubbling was heard coming from the pipes.
The exorcism must have freed some Sumerian ghost stuck in our MOPA's chilling pipes (we have strong reasons to believe it might have looked like this) because then the NPRO's radiator started getting cooler.
I also jiggled a bit with the valve while I was trying to unlock the safety nut, but I stopped when I noticed that the nut was stuck to the plastic support it is mounted on.
 
We're now watching the MOPA power's monitor to see if eventually all the tinkering succeeded.

 

[From Jenne:  When we first opened up the MOPA box, the NPRO's cooling fins were HOT.  This is a clear sign of something badbadbad.  They should be COLD to the touch (cooler than room temp).  After jiggling the needle valve, and hearing the water-rushing sounds, the NPRO radiator fins started getting cooler.  After ~10min or so, they were once again cool to the touch.  Good news.  It was a little worrisome however that just after our needle-valve machinations, the DTEC was going down (good), but the HTEMP started to rise again (bad).  It wasn't until after Alberto's tinkering that the HTEMP actually started to go down, and the power started to go up.  This is probably a lot to do with the fact that these temperature things have a fairly long time constant. 

Also, when we first went out to check on things, there was a lot more condensation on the water tubes/connections than I have seen before.  On the outside of the MOPA box, at the metal connectors where the water pipes are connected to the box, there was actually a little puddle, ~1cm diameter, of water. Steve didn't seem concerned, and we dried it off.  It's probably just more humid than usual today, but it might be something to check up on later.]

ELOG V3.1.3-